# NoisyGRPO: Incentivizing Multimodal CoT Reasoning via Noise Injection and Bayesian Estimation

### v1

@@ 3.1 Training Objective and Reward Function
Specifically, given a sample $(q,a,I)$ drawn from the dataset $\mathcal{D}$—where $q$ is the input query, $a$ is the ground-truth answer, and $I$ is the corresponding image—we sample $G$ rollouts $\{o_{i}^{n}\}_{i=1}^{G}$ using a noise-injected exploration policy $\pi^{\text{noise}}$, where each $o_{i}^{n}$ is sampled under a different noise level applied to the image. The corresponding group-wise advantage $\tilde{A}^{k}_{i}$ is estimated based on both trajectory reward and noise level. The training objective of NoisyGRPO is defined as follows:
@@ 3.2 Noise Injected Exploration Policy
To enhance the generalization of the multimodal RL framework, we introduce a noise-injected exploration strategy to construct our exploration policy $\pi^{\text{noise}}$, as shown in Figure 2. Different from injecting noise into the entire policy model as in [Igl2019GeneralizationIR], we adopt a targeted and structured strategy by applying multi-step Gaussian noise to the input image $I$ based on a diffusion-style forward process. For each rollout, we first sample a noise level $n_{i}\sim\mathcal{U}(0,1)$, representing a normalized noise step. We then perturb $I$ using the corresponding diffusion process at timestep $\lfloor n_{i}\cdot T\rfloor$, where $T$ is the total number of diffusion steps. $n_{i}=0$ yields the original clean image, while $n_{i}=1$ corresponds to the image at the final diffusion step, effectively resembling pure Gaussian noise. This produces a noisy image $I_{n_{i}}$ which is then used to generate the rollout $o_{i}^{n}$ under the exploration policy $\pi^{\text{noise}}$, resulting in a diverse set of responses $\{o_{i}^{n}\}_{i=1}^{G}$ conditioned on different noise intensities. Notably, noise injection is applied only during rollout collection to encourage exploration. Clean inputs are used during policy updates to ensure stable learning and preserve the accuracy of the learned policy $\pi_{\theta}$.
@@ 4.1 Training Data
To enhance the general Chain-of-Thought (CoT) reasoning ability of MLLMs through reinforcement learning, we adopt the visual question answering (VQA) portion of the MM-RLHF [MMRLHF] training set as our training data. This dataset spans diverse domains, including conversations, safety, multiple-choice questions, captions, and commonsense reasoning. MM-RLHF applies clustering and filtering techniques to curate a high-quality visual instruction-following dataset. In total, it contains 13k VQA samples: 1.2k yes/no questions, 1.3k multiple-choice questions, and 10k open-ended VQA questions. Notably, most of the training data lack structured answers; therefore, we use a text embedding model to compute the accuracy reward. While this may introduce some model bias, our Bayesian advantage estimation helps mitigate its impact.

### v3

@@ 3.1 Training Objective and Reward Function
Specifically, given a sample $(q,a,I)$ drawn from the dataset $\mathcal{D}$—where $q$ is the input query, $a$ is the ground-truth answer, and II is the corresponding image—we sample $G$ rollouts $\{o_{i}^{n}\}_{i=1}^{G}$ using a noise-injected exploration policy $\pi^{\text{noise}}$, where each $o_{i}^{n}$ is sampled under a different noise level applied to the image. The corresponding group-wise advantage $\tilde{A}^{k}_{i}$ is estimated based on both trajectory reward and noise level. The training objective of NoisyGRPO is defined as follows:
@@ 3.2 Noise Injected Exploration Policy
To enhance the generalization of the multimodal RL framework, we introduce a noise-injected exploration strategy to construct our exploration policy $\pi^{\text{noise}}$, as shown in Figure 2. Different from injecting noise into the entire policy model as in [Igl2019GeneralizationIR], we adopt a targeted and structured strategy by applying multi-step Gaussian noise to the input image II based on a diffusion-style forward process. For each rollout, we first sample a noise level $n_{i}\sim\mathcal{U}(0,1)$, representing a normalized noise step. We then perturb II using the corresponding diffusion process at timestep $\lfloor n_{i}\cdot T\rfloor$, where $T$ is the total number of diffusion steps. $n_{i}=0$ yields the original clean image, while $n_{i}=1$ corresponds to the image at the final diffusion step, effectively resembling pure Gaussian noise. This produces a noisy image $I_{n_{i}}$ which is then used to generate the rollout $o_{i}^{n}$ under the exploration policy $\pi^{\text{noise}}$, resulting in a diverse set of responses $\{o_{i}^{n}\}_{i=1}^{G}$ conditioned on different noise intensities. Notably, noise injection is applied only during rollout collection to encourage exploration. Clean inputs are used during policy updates to ensure stable learning and preserve the accuracy of the learned policy $\pi_{\theta}$.
@@ 4.1 Training Data
To enhance the general Chain-of-Thought (CoT) reasoning ability of MLLMs through reinforcement learning, we adopt the visual question answering (VQA) portion of the $M$-RLHF [MMRLHF] training set as our training data. This dataset spans diverse domains, including conversations, safety, multiple-choice questions, captions, and commonsense reasoning. $M$-RLHF applies clustering and filtering techniques to curate a high-quality visual instruction-following dataset. In total, it contains 13k VQA samples: 1.2k yes/no questions, 1.3k multiple-choice questions, and 10k open-ended VQA questions. Notably, most of the training data lack structured answers; therefore, we use a text embedding model to compute the accuracy reward. While this may introduce some model bias, our Bayesian advantage estimation helps mitigate its impact.